OpenStudy (anonymous):

What is the factored form of the expression? 4x^2 – 81y^2 A. (2x + 9)(2x – 9) B. (2x + 9y)(2x – 9y) C. (2x + 9y)2 D. (2x – 9y)2

Parth (parthkohli):

This is a difference of squares.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow a^2 - b^2 = (a + b)(a - b) \Longleftrightarrow (2x)^2 - (9y)^2 = (2x + 9y)(2x - 9y) }\)

Parth (parthkohli):

* \(\Large \color{MidnightBlue}{\Rightarrow (2x + 9y)(2x - 9y) }\)